Mira and Marie sell computers. Last month Mira sold 16 more computers than Marie. Together they sold 56. How many computers did Marie sell?

a. 20
b. 36
c. 16
d. 40
Answer #1

Let X = no of computers sold by mira

Y = No of computers sold by Marie

Given X = Y+16

X+Y = 56

Y+16 +Y = 56

2Y = 40

Y = 20

X+20 = 56

X = 36

No of computers sold by Marie = 20